### v3

@@ 3.3 GPT Score
The invoking of OpenAI APIs Tingiris and Kinsella (2021); Lappalainen and Narayanan (2023); Sun et al. (2023); Kublik and Saboo (2023) for ChatGPT services (e.g., GPT3.5, GPT4) allows automatic scoring of instruction tuning datasets. Recent studies on bringing LLMs as judges Zheng et al. (2024); Wang et al. (2023a); Zhu et al. (2023); Huang et al. (2024); Zeng et al. (2023); Chan et al. (2023) reveal that powerful language models like ChatGPT highly align with human preference on judging the quality of instructions and responses. Given a well-designed prompt with clear definition on grading criteria, ChatGPT produces justified quality scorings with explanations:
$GPTScore_{i}=G(x_{i},p_{G}),$ (18)
where $p_{G}$ denotes the prompt template that defines the task and grading scheme with format constraints on outputs. $G$ represents the quality score parsed from the GPT response. Samples with high $GPTScore_{i}$ can be selected using Eqs. 4 and  5.
Prompt $p_{G}$ for scoring $x_{i}$ with instruction (input) and response in the <dimension> We would like to request your feedback on the performance of AI assistant in response to the instruction and the given input displayed following. Instruction: <instruction> Input: <input> Response: <response> Please rate according to the <dimension> of the response to the instruction and the input. Each assistant receives a score on a scale of 0 to 5, where a higher score indicates higher level of the <dimension>. Please first output a single line containing the value indicating the scores. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
@@
Chen et al. (2023b) proposes a surprisingly easy-yet-effective method that directly uses GPT3.5 to score datapoints in terms of helpfulness and accuracy (see the detailed prompt 3.3). Both instructions and responses are scored on a scale from 00 to $5$ and experimental results show that general instruction datasets, except coding-related samples, can be distilled into smaller subsets for better downstream performance. Bukharin and Zhao (2023) follows Chen et al. (2023b) for filtering Alpaca Taori et al. (2023). Chen and Mueller (2024) employs the BSDetector Chen and Mueller (2023) to estimate the confidence of GPT3.5/GPT4 on the give instruction-response pair. It takes both the self-consistency and direct scoring into consideration. Only highly confident samples are kept for fine-tuning domain-specific LLMs and those less confident ones are corrected automatically by these LLMs. Xu et al. (2023b) directly evaluates instruction datasets in terms of accuracy, explanation, clarity, and difficulty for weighted scorings from GPT4. Then, both hand-crafted indicators (i.e., lengthwise semantic evaluation) and GPT4 scorings are employed for final ranking. Liu et al. (2023b) argues that the direct scoring of GPT4 on one single instruction sample is not well-calibrated and instead gives relative ranking of multiple instruction variants at once. The complexity of instructions Xu et al. (2023a) and the quality of instruction-response pairs are sequentially obtained from GPT3.5. Zhang et al. (2024c) uses GPT scorings to judge: 1) whether the given text contains mathematical contents; 2) and if yes, whether these maths contents are of high quality for education purpose. Such scores are proved more effective than traditional “mathematical" classifiers Paster et al. (2023). Lu et al. (2023a) proposes to use ChatGPT for annotating open-ended, fine-grained intention tags on open datasets. Then, the quality of the tag dataset is evaluated by humans and GPT4 in terms of tagging precision and consistency. Instead of fully relying on the GPT4, Li et al. (2023c) exploits the model under investigation itself (e.g., LLaMA 65B) to iteratively derive quality scores on each augmented example on a 5-point scale. Then a curated clean set is chosen via Eq. 4.
QuRator Wettig et al. (2024) manually defines quality criterion such as writing style, facts and trivia, educational value, and required expertise. Then, quality comparison is conducted on two instruction-response samples via GPT3.5 scoring. Such pairwise scorings are used to fine-tune a sheared-LLaMA 1.3B model Xia et al. (2023) in a manner similar to DPO Ouyang et al. (2022); Rafailov et al. (2024). It is noted that pairwise scoring Ouyang et al. (2022); Dubois et al. (2024); Zeng et al. (2023); Liu et al. (2023b) have been found more reliable, consistent, and unbiased than individual scoring Gunasekar et al. (2023); Chen et al. (2023b) during GPT-based quality analysis.
@@
Closed-source LLMs such as ChatGPT enjoy a high level of alignment with human preference and therefore can be utilized to score data quality. It would be more cost-efficient to collect few (e.g., <100K) GPT-scored samples first and then fine-tune an open-source LLM for quality measurement on massive instruction corpus.
@@ 3.4 Human Evaluation
Human annotation and evaluation is indispensable in constructing preference alignment datasets Wang et al. (2023b); Ouyang et al. (2022) for helpfulness, honesty, and harmlessness. Specifically, human annotators deliver grading results following specific criteria in multiple dimensions:
@@
Guidelines (excerpts) for human annotations # Guidelines Below is a list of guidelines that should be adhered to for each possible task available when building the dataset. To see some examples of how the guidelines can be applied, visit the examples document. ## 1. General rules - Always make sure to read and understand the guidelines to each task before fulfilling it. - Try to follow the guidelines as closely as possible. - If you are unsure whether a message violates a guidelines, contact us at our Discord. - Use the thumbs-up/thumbs-down system to further mark messages that are of high or low quality. ## 2. Providing an assistant reply #assistant-reply ### Do: - Remain polite and treat the user with respect, even when not given the same courtesy.
@@
The OpenAssistant Köpf et al. (2024) dataset is featured by its high-quality human-generated, human-annotated multi-lingual conversations for both instruction tuning and reinforcement learning from human feedback (see the guidelines excerpts 3.4). For each instruction-response pair along the conversation tree, the human annotators are asked to categorize them according to three dimensions: spam detection, guideline adherence, and quality. The quality score is rated on a five-point Likert scale across aspects including quality, creativity, humorousness, politeness, and harmlessness. These scores are used to sort instructions for analysis and preference optimization of LLMs.  Lu et al. (2023a) enrolls human annotators to provide judgements on the tagging of each instruction. To verify the quality scores provided by humans, counterfactual cases are prepared respectively for precision and consistency tasks. Results show that human annotators have low false positive rates at tagging precision, but lack proof of confidence on their original quality judgements. Zhou et al. (2024a) proposes to use human annotators for creation of small-yet-effective instruction datasets. To collect questions and answers from various sources, simple hand-crafted indicators such as text length are used to filter low-quality datapoints. Then, high quality instruction-response pairs are manually selected (750) and written (250) via subjective quality control. The databricks-dolly dataset Conover et al. (2023) contains 15K human-generated instruction-response pairs. Although quality is emphasized during large-scale annotation, imperfect samples still exist where low-quality and inaccurate responses, incomplete and vague instructions, problematic texts with toxic language and grammar errors are found He et al. (2024).
@@
Human evaluation play a irreplaceable role in quality control of preference alignment. To reduce the inter-annotator inconsistency, detailed guidelines should be prepared for quality measurement. In addition, supplementary quality measures such as GPT-Scores can be provided for manually evaluating and selecting high-quality datasets.
@@ 4 Diversity-based Selection
In this section, we introduce methods that emphasize the diversity of instruction datasets. When it comes to diversity, existing researches either measure the individual diversity of each sample (e.g., lexical and semantic richness) or the overall diversity of the entire dataset (e.g., the volume of the enclosed embedding space). Instruction datapoints whose tasks and domains are of minority classes in a long-tailed distribution are preferred during subset selection. Such sampling philosophy strikes to maintain or approximate the spread of the original embedding clusters but with much less sparsity.
@@ 4.1 Hand-crafted Indicators
The diversity of datasets is the key to develop less biased, more generalizable machine learning models. However, recent studies Zhao et al. (2024c, b) show that existing vision and language datasets do not share a unified and concrete definition of diversity in terms of dataset composition, source, domain, subject, annotator, and promote (fairness). With respect to the diversity measures specific in instruction tuning datasets, hand-crafted indicators, similar to Eq. 3 in traditional NLP studies, can be used as a good starting point.
@@
where $Unique(x_{i})$ denotes the set of unique tokens present in $x_{i}$. To reduce the sensitivity of TTR to the variation of text length, several studies Covington and McFall (2008, 2010); Kettunen (2014); Matlach et al. (2021) standardized the length by introducing logarithms or n-grams into the formula.
Later, computational approaches to measure lexical diversity have been developed such as vocabulary diversity (vocd-D) Malvern and Richards (1997); Malvern et al. (2004); Silverman and Ratner (2002); deBoer (2014), the measure of textual lexical diversity (MTLD) McCarthy and Jarvis (2010); Jarvis and Daller (2013), and hypergeometric distribution diversity (HD-D) Jarvis (2013); McCarthy (2005). All these metrics require multi-step computation for approximation. Specifically for vocd-D, random sampling is first performed on $x_{i}$ for a series of sub-sequences with varying lengths $k$ (e.g., 10, 20, 30 tokens). Then, $TTR^{k}$ is:
@@
The HD-D shares the same idea behind vocd-D but stems from the hypergeometric distribution McCarthy and Jarvis (2010). With $M$-times sampling, the HD-D represents the probability of drawing a certain number of tokens of the given type from the subsequence of $x_{i}$ with a particular size $k$:
$HD\text{-}D_{i}=$ $\sum_{t=1}^{|Unique(x_{i})|}\frac{1}{M}\sum^{M}_{m=1}\mathbbm{1}(x_{i(n)}^{m}=u_{t},$ (25)
$\exists 1\leq n\leq|x_{i}^{m}|),u_{t}\in Unique(x_{i}),$
$x_{i}^{m}=$ $x_{i(j\leq,<j+k)},\forall k>0,\forall 1\leq j\leq|x_{i}|-k.$
Other variants of TTR indicators such as MTTRSS Malvern et al. (2004), MSTTR Malvern et al. (2004), MATTR Covington and McFall (2010), and MTLD-W Vidal and Jarvis (2020); Kyle et al. (2021) all target at improving the solutions to two fundamental problems Bestgen (2023): 1) the sensitivity of indicators to text length, and 2) the impact of the indicator parameters. Li et al. (2015) proposes two rather simplified TTR scores as $distinct\text{-}1$ and $distinct\text{-}2$, where the number of distinct unigrams and bigrams of $x_{i}$ are respectively devided by the total number of words. Many other studies Cao and Clark (2017); Zhu et al. (2018); Shu et al. (2019); Tevet and Berant (2020) extend the application of n-gram-based diversity measures for model-generated responses.
Apart from lexical diversity, there exists many efficient diversity indicators that are built upon the semantics of each example. Dong et al. (2011) proposes to approximate k-nearest neighbor (k-NN) graph Peterson (2009) with arbitrary similarity measures on semantic embeddings of large-scale datasets. Such efficient construction of a k-NN graph allows the distance of $x_{i}$ to its $j$-th nearest neighbors to be a feasible diversity measure:
@@
where $N_{j}(x_{i})$ denotes the $j$-th closest neighbor of $x_{i}$ in the embedding space projected by $g(\cdot)$. The common choices of the distance function $d(\cdot,\cdot)$ include the Euclidean distance, cosine distance, and Jaccard coefficient distance Huang et al. (2008). The projection from text (e.g., instruction-response pairs) into the embedding space can be achieved with pre-trained sentence BERT Reimers and Gurevych (2019); Feng et al. (2020), where an additional pooling operation is performed on the final output of BERT Devlin et al. (2018) for sentence embeddings. Note that a higher $kNN_{i}$ implies that the sample $x_{i}$ is more unique and should be kept in subset selection for higher diversity. Due to the fine-grained representation capability of BERT, existing hand-crafted indicators often rely on BERT embeddings for similarity or diversity measurement Tevet and Berant (2020); Zhang et al. (2019); Larson et al. (2019); Yauney et al. (2023).
To improve the generalization of diversity measure, Xu et al. (2023b) argues that the statistics of feature embedding of each sample itself should be considered. It does not require additional prior knowledge on the structure of embeddings. Given all datapoints $x_{i}\in S$, their semantic embeddings from any sentence encoder can be represented as $X=[g(x_{1}),g(x_{2}),...,g(x_{N})]\in\mathcal{R}^{|S|\times H}$. The row variance $Var_{i}$ of each embedding $g(x_{i})$ in the reduced dimensional space $\mathcal{R}^{|S|\times k}$ by principal components analysis (PCA) Wold et al. (1987) is used as the diversity indicator:
@@
When it comes to the overall diversity of a dataset $S$, the average distance of any sample $x_{i}$ to its closest neighbor in the dataset, namely $kNN_{i}$, can be leveraged intuitively:
@@
Such a diversity measure has been widely used in dataset construction and content retrieval Stasaski et al. (2020); Stasaski and Hearst (2022); Mithun et al. (2019); Spyromitros-Xioufis et al. (2015); Sun et al. (2024a); Ionescu et al. (2018). Du and Black (2019) simply performs clustering on all samples with k-means Ikotun et al. (2023) into $K$ clusters ($C_{1}$,$C_{2}$,…,$C_{K}$) in the embedding space, and then uses the cluster inertia as diversity indicators:
@@
Lai et al. (2020) develops a diversity metric on the dispersion of a cluster induced by embeddings of all samples, where the cluster is approximated by a multi-variate Gaussian distribution:
@@
where $H$ is the dimension of the projected embeddings $g(x_{i})\in\mathcal{R}^{H}$ and $\sigma_{j}$ denotes the radius of the ellipsoid along the $j$-th axis of the dataset $S$. The inter-cluster (class) distance can also be used for diversity measure Dang and Verma (2024):

### v4

@@ 3.3 GPT Score
The invoking of OpenAI APIs Tingiris & Kinsella (2021); Lappalainen & Narayanan (2023); Sun et al. (2023); Kublik & Saboo (2023) for ChatGPT services (e.g., GPT3.5, GPT4) allows automatic scoring of instruction tuning datasets. Recent studies on bringing LLMs as judges Zheng et al. (2024); Wang et al. (2023a); Zhu et al. (2023); Huang et al. (2024); Zeng et al. (2023); Chan et al. (2023) reveal that powerful language models like ChatGPT highly align with human preference on judging the quality of instructions and responses. Given a well-designed prompt with clear definition on grading criteria, ChatGPT produces justified quality scorings with explanations for the raw textual instruction data $I_{i}$:
$GPTScore_{i}=G(\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I_{i}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},p_{G}),$ (21)
where $p_{G}$ denotes the prompt template that defines the task and grading scheme with format constraints on outputs (see Fig. 4). The $G(\cdot,\cdot)$ represents the process of quality scoring and response parsing, which produces the GPT score $GPTScore_{i}$. Samples with high $GPTScore_{i}$ can be selected using Eqs. 7 and  8.
Prompt $p_{G}$ for scoring $I_{i}$ with instruction (input) and response in the <dimension> We would like to request your feedback on the performance of AI assistant in response to the instruction and the given input displayed following.
Instruction: <instruction>
Input: <input>
Response: <response>
Please rate according to the <dimension> of the response to the instruction and the input. Each assistant receives a score on a scale of 0 to 5, where a higher score indicates higher level of the <dimension>.
Please first output a single line containing the value indicating the scores. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
Figure 4: The prompt $p_{G}$ for scoring the raw text $I_{i}$ with ChatGPT.
@@
Chen et al. (2023b) propose a surprisingly easy-yet-effective method that directly uses GPT3.5 to score datapoints in terms of helpfulness and accuracy. Both instructions and responses are scored on a scale from 00 to $5$ and experimental results show that general instruction datasets, except coding-related samples, can be distilled into smaller subsets for better downstream performance. Bukharin & Zhao (2023) follow Chen et al. (2023b) for filtering Alpaca Taori et al. (2023). Chen & Mueller (2024) employ the BSDetector Chen & Mueller (2023) to estimate the confidence of GPT3.5/GPT4 on the give instruction-response pair. It takes both the self-consistency and direct scoring into consideration. Only highly confident samples are kept for fine-tuning domain-specific LLMs and those less confident ones are corrected automatically by these LLMs. Xu et al. (2023b) directly evaluate instruction datasets in terms of accuracy, explanation, clarity, and difficulty for weighted scorings from GPT4. Then, both hand-crafted indicators (i.e., lengthwise semantic evaluation) and GPT4 scorings are employed for final ranking. Liu et al. (2023b) argue that the direct scoring of GPT4 on one single instruction sample is not well-calibrated and instead gives relative ranking of multiple instruction variants at once. The complexity of instructions Xu et al. (2023a) and the quality of instruction-response pairs are sequentially obtained from GPT3.5. Zhang et al. (2024c) use GPT scorings to judge: 1) whether the given text contains mathematical contents; 2) and if yes, whether these maths contents are of high quality for education purpose. Such scores are proved more effective than traditional "mathematical" classifiers Paster et al. (2023). Lu et al. (2023a) propose to use ChatGPT for annotating open-ended, fine-grained intention tags on open datasets. Then, the quality of the tag dataset is evaluated by humans and GPT4 in terms of tagging precision and consistency. Instead of fully relying on the GPT4, Li et al. (2023c) exploit the model under investigation itself (e.g., LLaMA 65B) to iteratively derive quality scores on each augmented example on a 5-point scale. Then a curated clean set is chosen via the Eq. 7. QuRator Wettig et al. (2024) manually define quality criterion such as writing style, facts and trivia, educational value, and required expertise. Then, quality comparison is conducted on two instruction-response samples via GPT3.5 scoring. Such pairwise scorings are used to fine-tune a sheared-LLaMA 1.3B model Xia et al. (2023) in a manner similar to DPO Ouyang et al. (2022); Rafailov et al. (2024). It is noted that the pairwise scoring Ouyang et al. (2022); Dubois et al. (2024); Zeng et al. (2023); Liu et al. (2023b) has been found more reliable, consistent, and unbiased than the individual scoring Gunasekar et al. (2023); Chen et al. (2023b) during GPT-based quality analysis.
@@
Closed-source LLMs such as ChatGPT enjoy a high level of alignment with human preference and therefore can be utilized for quality scoring. It would be cost-efficient to collect few (e.g., <100K) GPT-scored samples first and then fine-tune an open-source LLM for quality measurement on massive corpus.
@@ 3.4 Human Evaluation
Evaluation with human-in-the-loop is indispensable in constructing preference alignment datasets Wang et al. (2023b); Ouyang et al. (2022) for helpfulness, honesty, and harmlessness. Specifically, human annotators deliver gradings following specific criteria (see Fig. 5) in multiple dimensions:
@@
Guidelines (excerpts) for human annotations # Guidelines
Below is a list of guidelines that should be adhered to for each possible task available when building the dataset. To see some examples of how the guidelines can be applied, visit the examples document.
## 1. General rules
- Always make sure to read and understand the guidelines to each task before fulfilling it. - Try to follow the guidelines as closely as possible. - If you are unsure whether a message violates a guidelines, contact us at our Discord.
- Use the thumbs-up/thumbs-down system to further mark messages that are of high or low quality.
## 2. Providing an assistant reply #assistant-reply
### Do:
- Remain polite and treat the user with respect, even when not given the same courtesy.
Figure 5: The guidelines (thumbnails) for human experts to create and annotate instruction datasets.
@@
The OpenAssistant Köpf et al. (2024) dataset is featured by its high-quality human-generated, human-annotated multi-lingual conversations for both instruction tuning and reinforcement learning from human feedback. For each instruction-response pair along the conversation tree, the human annotators are asked to categorize them according to three dimensions: spam detection, guideline adherence, and quality. The quality score is rated on a five-point Likert scale across aspects including quality, creativity, humorousness, politeness, and harmlessness. These scores are used to sort instructions for analysis and preference optimization of LLMs.  Lu et al. (2023a) enroll human annotators to provide judgements on the tagging of each instruction. To verify the quality scores provided by humans, counterfactual cases are prepared respectively for precision and consistency tasks. Results show that human annotators have low false positive rates at tagging precision, but lack proof of confidence on their original quality judgements. Zhou et al. (2024a) propose to use human annotators for creation of small-yet-effective instruction datasets. To collect questions and answers from various sources, simple hand-crafted indicators such as text length are used to filter low-quality datapoints. Then, high quality instruction-response pairs are manually selected (750) and written (250) via subjective quality control. The databricks-dolly dataset Conover et al. (2023) contains 15K human-generated instruction-response pairs. Although quality is emphasized during large-scale annotation, imperfect samples still exist. For example, low-quality and inaccurate responses, incomplete and vague instructions, problematic texts with toxic language and grammar errors are found He et al. (2024).
@@
Human evaluation plays an irreplaceable role in quality control of preference alignment. To reduce the inter-annotator inconsistency, detailed guidelines should be prepared for quality measurement. In addition, supplementary quality measures such as GPT scores can be provided for reference during evaluation and selection of high-quality datapoints.
@@ 4 Diversity-based Selection
In this section, we introduce methods that emphasize the diversity of instruction datasets. When it comes to diversity, existing researches either measure the individual diversity of each sample (e.g., lexical and semantic richness) or the overall diversity of the entire dataset (e.g., the volume of the enclosed embedding space). Datapoints whose tasks and domains are of minority classes in a long-tailed distribution are preferred during subset selection. Such sampling philosophy strikes to maintain or approximate the spread of the original embedding clusters but with much less sparsity. In the case of diversity, the evaluation function $q(x_{i})$ can be represented in the unified formulation as:
$q(x_{i})=f_{d}(q_{L}(x_{i}),q_{S}(x_{i})),$ (23)
where $q_{L}$ measures the lexical diversity of $x_{i}$ and $q_{S}$ assesses the semantic diversity. The $f_{d}$ denotes the aggregation function in diversity measurement. Typically, $q_{L}$ often investigates the diversity of n-grams, tokens, words, and sequences. Complementarily, $q_{S}$ emphasizes semantic diversity that the variety of representations of the selected datapoints should be maximized in the embedding space. Both the two aspects of diversity can be sequentially or jointly considered to remove any duplicates in the instruction datasets.
@@ 4.1 Hand-crafted Indicators
The diversity of datasets is the key to develop less biased, more generalizable machine learning models. However, recent studies Zhao et al. (2024c; b) show that existing datasets do not share a unified and concrete definition of diversity in terms of dataset composition, source, domain, subject, annotator, and promote (fairness). With respect to the diversity measures specific in instruction tuning datasets, hand-crafted indicators, similar to Eq. 6 in traditional NLP studies, can be used as a good starting point.
@@
where $Unique(x_{i})$ denotes the set of unique tokens present in $x_{i}$. To reduce the sensitivity of TTR to the variation of text length, several studies Covington & McFall (2008; 2010); Kettunen (2014); Matlach et al. (2021) standardized the length by introducing logarithms or n-grams into the formula.
Later, computational approaches to measure lexical diversity have been developed such as vocabulary diversity (vocd-D) Malvern & Richards (1997); Malvern et al. (2004); Silverman & Ratner (2002); deBoer (2014), the measure of textual lexical diversity (MTLD) McCarthy & Jarvis (2010); Jarvis & Daller (2013), and hypergeometric distribution diversity (HD-D) Jarvis (2013); McCarthy (2005). All these metrics require multi-step computation for approximation. Specifically for vocd-D, random sampling is first performed on $x_{i}$ for a series of sub-sequences with varying lengths $k$ (e.g., 10, 20, 30 tokens). Then, $TTR^{k}$ is:
@@
The HD-D shares the same idea behind vocd-D but stems from the hypergeometric distribution McCarthy & Jarvis (2010). With $M$-times sampling, the HD-D represents the probability of drawing a certain number of tokens of the given type from the subsequence of $x_{i}$ with a particular size $k$:
$HD\text{-}D_{i}=$ $\sum_{t=1}^{|Unique(x_{i})|}\frac{1}{M}\sum^{M}_{m=1}\mathbbm{1}(x_{i(n)}^{m}=u_{t}),1\leq n\leq|x_{i}^{m}|,$ (29)
$u_{t}\in Unique(x_{i}),x_{i}^{m}=x_{i(j\leq,<j+k)},\forall k>0,\forall 1\leq j\leq|x_{i}|-k.$
Variants of TTR indicators such as MTTRSS Malvern et al. (2004), MSTTR Malvern et al. (2004), MATTR Covington & McFall (2010), and MTLD-W Vidal & Jarvis (2020); Kyle et al. (2021) all target at the solutions to two fundamental problems Bestgen (2023): 1) the sensitivity of indicators to text length, and 2) the impact of the indicator parameters. Li et al. (2015) propose two rather simplified TTR scores as $distinct\text{-}1$ and $distinct\text{-}2$, where the number of distinct unigrams and bigrams of $x_{i}$ are respectively divided by the total number of tokens. Other studies Cao & Clark (2017); Zhu et al. (2018); Shu et al. (2019); Tevet & Berant (2020) extend the application of n-gram-based diversity for model-generated responses.
Apart from lexical diversity, there exists many efficient diversity indicators that are built upon the semantics of each example. Dong et al. (2011) propose to approximate k-nearest neighbor (k-NN) graph Peterson (2009) with arbitrary similarity measures on semantic embeddings of large-scale datasets. Such efficient construction of a k-NN graph allows the distance of $x_{i}$ to its $j$-th nearest neighbors to be a feasible diversity measure:
@@
where $N_{j}(x_{i})$ denotes the $j$-th closest neighbor of $x_{i}$ in the embedding space projected by $g(\cdot)$. The common choices of the distance function $d(\cdot,\cdot)$ include the Euclidean distance, cosine distance, and Jaccard coefficient distance Huang et al. (2008). The projection from text (e.g., instruction-response pairs) into the embedding space can be achieved with pre-trained sentence BERT Reimers & Gurevych (2019); Feng et al. (2020), where an additional pooling operation is performed on the final output of BERT Devlin et al. (2018) for sentence embeddings. Note that a higher $kNN_{i}$ implies that the sample $x_{i}$ is more unique and should be kept in subset selection for higher diversity. Due to the fine-grained representation capability of BERT, existing hand-crafted indicators often rely on BERT embeddings for similarity or diversity measurement Tevet & Berant (2020); Zhang et al. (2019); Larson et al. (2019); Yauney et al. (2023).
To improve the generalization of diversity measure, Xu et al. (2023b) argue that the statistics of feature embedding of each sample itself should be considered. It does not require additional prior knowledge on the structure of embeddings. Given all datapoints $x_{i}\in S$, their semantic embeddings from any sentence encoder can be represented as $X=[g(x_{1}),g(x_{2}),...,g(x_{N})]\in\mathcal{R}^{|S|\times H}$. The row variance $Var_{i}$ of each embedding $g(x_{i})$ in the reduced dimensional space $\mathcal{R}^{|S|\times k}$ by principal components analysis (PCA) Wold et al. (1987) is used as the diversity indicator:
@@
When it comes to the overall diversity of a dataset $S$, the average distance of any datapoint $x_{i}$ to its closest neighbor in the dataset, namely $kNN_{i}$, can be leveraged intuitively:
@@
Such a diversity measure has been widely used in dataset construction and retrieval Stasaski et al. (2020); Stasaski & Hearst (2022); Mithun et al. (2019); Spyromitros-Xioufis et al. (2015); Sun et al. (2024a); Ionescu et al. (2018). Du & Black (2019) simply perform clustering on all samples with k-means Ikotun et al. (2023) into $K$ clusters ($C_{1}$,$C_{2}$,…,$C_{K}$) in the embedding space, and then uses the cluster inertia as diversity indicators:
@@
Lai et al. (2020) develop a diversity metric on the dispersion of a cluster induced by embeddings of all samples, where the cluster is approximated by a multi-variate Gaussian distribution:
@@
where $H$ is the dimension of the projected embeddings $g(x_{i})\in\mathcal{R}^{H}$ and $\sigma_{j}$ denotes the radius of the ellipsoid along the $j$-th axis of the dataset $S$. The inter-cluster (class) distance can also be used for diversity measure Dang & Verma (2024):